OpenStudy (anonymous):

How do you simplify the cube root of 32?

OpenStudy (anonymous):

Make the prime factors of 32.. \[32 = 2 \times 2 \times 2 \times 2 \times 2\] Now from cube roots, if we have 3 common digits, we can bring that out of cube root. Here, combine three 2's together and underline them: \[\large \sqrt[3]{32} = \sqrt[3]{\underline{2 \times 2 \times 2} \times 2 \times 2}\] Now you can bring one 2 for that three 2's leaving behind two 2's inside like: \[\large \sqrt[3]{32} = 3 \sqrt[3]{2 \times 2}\]

OpenStudy (anonymous):

why is the number to the left of the radical a 3?

OpenStudy (anonymous):

because we took out three numbers?

OpenStudy (anonymous):

Here you cannot bring more out of the cube root brackets, because there are only two 2's inside, if it would have three 2's inside then you can bring one out of that three also..

OpenStudy (anonymous):

That is representing cube root, root means brackets, and cube means 3 that is what I have written there: Cube root of a number is shown like this: \[\huge \color{blue}{\sqrt[3]{Number}}\]

OpenStudy (anonymous):

so the answer is 3*3rt(4)?

OpenStudy (anonymous):

said as "Three times the cube root of 4"

OpenStudy (anonymous):

Yep but write it in proper way like: